Scores on a scholarship aptitude examination for senior high school student in a top university in the Visayas are normally distributed with a mean of 70 and standard deviation of 10. What is the highest score that will place an applicant in the bottom 30% (30th percentile) of the distribution?


1
Expert's answer
2021-01-19T04:29:57-0500

"Score_{0.30} = 70 + 10*Z_{0.30}=70 + 10*(-0.5244) = 64.7560\\"